Given the following information:
P(A) = 0.59
P(B) = 0.5
P(A and B) = 0.39
P(B given A) = 0.2
What is the probability of A or B? Round your answer to the nearest hundredth as needed.
P
(A or B) =

Answer by ikleyn(52939)

Given the following information:
P(A) = 0.59
P(B) = 0.5
P(A and B) = 0.39
P(B given A) = 0.2
What is the probability of A or B? Round your answer to the nearest hundredth as needed.
P(A or B) =
~~~~~~~~~~~~~~~~~~

    P(A or B) = P(A) + P(B) - P(A and B) = 0.59 + 0.5 - 0.39 = 0.7.    ANSWER

Solved and explained.